OpenStudy (anonymous):

Is - sqrt ( 54) rational or not ?

OpenStudy (anonymous):

\[ -\sqrt{54} = -\sqrt{2\cdot 3^3} = -3\sqrt{2}\sqrt{3} \]It doesn't look rational.